t (t) = + lg(t) = 2t
for the following exercises, eliminate the parameter t to rewrite the parametric equation as a Cartesian equation
Answer #1

xct) - 19. t , Чt) = 2 t дуо 9 = 2 t: 81 2. -) х = (81) (91-) 9) 8°/- 47, = 8 (924) 8 Я 3 48 8 х = 8 (94) А CS Scanned with Ca